OpenStudy (anonymous):

Tickets for a school play cost $4 for adults and $2 for students. At the end of the play, the school sold a total of 105 tickets and collected $360. Question 4 Unsaved Write a linear system. Let x be the number of adult tickets sold and let y be the number of student tickets sold.

OpenStudy (anonymous):

I came up with this 105/(4x+2y)=360

OpenStudy (anonymous):

Ah sorry I relazed what I did wrong. I'm just writing an equation not a linear system it should be 4x+2y=$360 x+y=105

OpenStudy (anonymous):

Thanks Bangokokgarrett

OpenStudy (anonymous):

good job!